OpenStudy (anonymous):

A navigator on a ship sights a lighthouse at a bearing of N36°E. After traveling 6.0 miles at a heading of 332°, the ship sights the lighthouse at a bearing of S82°E. How far is the ship from the lighthouse at the second sighting?

OpenStudy (anonymous):

6.1 miles

OpenStudy (anonymous):

What is the solution?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\theta = 360 -332 = 28° \]

OpenStudy (anonymous):

C = 82 - 28 =54°

OpenStudy (anonymous):

A = 36 + 28 = 64°

OpenStudy (anonymous):

B= 180 - 54 - 64 = 62°

OpenStudy (anonymous):

Then use law of sines.

OpenStudy (anonymous):

\[\frac{ 8 }{ \sin62 } = \frac{ a }{ \sin64 } \]

OpenStudy (anonymous):

oh my mistake. 6/sin 62